# The spectrum of Dirichlet-to-Neumann maps for radial conductivities

### v1

@@ 3.5. Explicit examples
We conclude this section by presenting a family of conductivities for which the Born approximation can be computed explicitly.
@@
    The presence of the zero resonance for $d\geq 3$ results in a Born approximation $\gamma^{\mathrm{B}}$ with a logarithmic singularity at the origin (and $V^{\mathrm{B}}$ with a $|x|^{-2}$ singularity), as predicted by Proposition˜3.10 and ˜5.
@@ 4.1. Hölder stability
provided that $\gamma_{1},\gamma_{2}\in\mathcal{K}\cap W^{2,p}(\mathbb{B}^{d})$, for some $p>d/2$, and that $\gamma_{1}|_{\mathbb{S}_{d-1}}=\gamma_{2}|_{\mathbb{S}_{d-1}}$. Indeed, by ˜1 and (2.11) we have
$|\lambda_{k}(\gamma_{1})-\lambda_{k}$ $(\gamma_{2})|\leq\left|\sigma_{k}[\Delta(\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2})]\right|+\left|\partial_{\nu}\gamma_{1}|_{\mathbb{S}_{d-1}}-\partial_{\nu}\gamma_{2}|_{\mathbb{S}_{d-1}}\right|$
$=\left|\sigma_{k}[\Delta(\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2})]\right|+\left|\sigma_{0}[\Delta(\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2})]\right|$
$\leq 2\sup_{k\in\mathbb{N}}\left|\sigma_{k}[\Delta(\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2})]\right|\leq 2\left\lVert\Delta(\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2})\right\rVert_{L^{1}(\mathbb{B}^{d})}\leq 2\left\lVert\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2}\right\rVert_{W^{2,1}(\mathbb{B}^{d})}.$
It is well known that in general $\omega_{\mathcal{K}}$ in (4.1) cannot be taken better than logarithmic when $\mathcal{K}$ includes a ball in a Sobolev space, [Man01, KRS21]. ˜4 improves estimate (4.2) to a Hölder modulus of continuity, and gets rid of the assumption $\gamma_{1}|_{\mathbb{S}_{d-1}}=\gamma_{2}|_{\mathbb{S}_{d-1}}$.
@@ 4.2. Proof of ˜4
then, if $a_{j}:=\gamma_{j}|_{\mathbb{S}^{d-1}}$, $b_{j}:=\partial_{\nu}\gamma_{j}|_{\mathbb{S}^{d-1}}$ one has the local estimate
@@
(4.5) $\left\lVert V_{1}-V_{2}\right\rVert_{L^{1}(U_{s})}<C_{d,N,p,s}\,\left\lVert V^{\mathrm{B}}_{1}-V^{\mathrm{B}}_{2}\right\rVert_{L^{1}(U_{s})}^{\frac{p}{2p-1}}.$
@@
Let $d\geq 2$, $N>0$ and $d/2<p<\infty$. There exists a constant $\beta(d,p)$ such that if $\left\lVert V\right\rVert_{L^{p}(\mathbb{B}^{d})}<N$, then $\left\lVert V^{\mathrm{B}}\right\rVert\leq\beta(d,p)N$.
@@
Since the conductivities are radial and (1.5) holds, by Lemma˜2.3 ii) we have for some $C>0$ that
@@
where $a_{j}:=\gamma_{j}|_{\mathbb{S}^{d-1}}$, $b_{j}:=\partial_{\nu}\gamma_{j}|_{\mathbb{S}^{d-1}}$. Locally, the same bound holds:
@@
Conditions (1.10) on $\gamma_{j}$, $j=1,2$ imply a corresponding bound for $\left\lVert V_{j}\right\rVert_{L^{p}}$ using that $\Delta\gamma_{j}=V_{j}\sqrt{\gamma_{j}}$, so the first estimates in (4.4) and (4.6) hold. We now consider the global and local cases separately.
Proof of local estimate i). First we claim that for any given $0<\varepsilon<1$, there exists a $0<\delta<1$ (depending on $\varepsilon,d,K,N,s$) such that $\left\lVert V^{\mathrm{B}}_{1}-V^{\mathrm{B}}_{2}\right\rVert_{L^{1}(U_{s})}<\varepsilon$ holds if $\left\lVert\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2}\right\rVert_{W^{2,1}(\mathbb{B}^{d})}<\delta$ holds for $\delta$. This can be proved using (2.19) together with $K^{-1}\leq\gamma_{j}(x)\leq K$ which imply that
$\left\lVert V^{\mathrm{B}}_{1}-V^{\mathrm{B}}_{2}\right\rVert_{L^{1}(U_{s})}=\left\lVert(2a_{1})^{-1}(2a_{1}V^{\mathrm{B}}_{1}-2a_{2}V^{\mathrm{B}}_{2})+a_{2}V^{\mathrm{B}}_{2}(a_{1}^{-1}-a_{2}^{-1})\right\rVert_{L^{1}(U_{s})}\\ \leq\frac{K}{2}\left\lVert\Delta\gamma^{\mathrm{B}}_{1}-\Delta\gamma^{\mathrm{B}}_{2}\right\rVert_{L^{1}(U_{s})}+|a_{1}^{-1}-a_{2}^{-1}||a_{2}|\left\lVert V^{\mathrm{B}}_{2}\right\rVert_{L^{1}(U_{s})},\$
we now use that $|a_{1}^{-1}-a_{2}^{-1}|\leq C_{K}|a_{1}-a_{2}|$, since $K^{-1}\leq a_{j}\leq K$, Lemma˜4.4 to bound the norm of $V^{\mathrm{B}}_{2}$, and (4.9), to obtain
@@
$\left\lVert\gamma_{1}-\gamma_{2}\right\rVert_{W^{2,1}(U_{s})}<C_{N,K,d,p,s}\,\left\lVert\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2}\right\rVert_{W^{2,1}(U_{s})}^{\frac{p}{2p-1}}+|a_{1}-a_{2}|+|b_{1}-b_{2}|.$
@@
Proof of global estimate ii) Using (1.10) and (4.8), the same arguments of the local case can be used to prove that
(4.11) $\left\lVert V^{\mathrm{B}}_{1}-V^{\mathrm{B}}_{2}\right\rVert_{L^{1}(\mathbb{B}^{d})}\leq C(d,K,N)\left\lVert\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2}\right\rVert_{W^{2,1}(\mathbb{B}^{d})}.$
Therefore, there exits a $0<\delta<1$ such that (4.7) holds if $\left\lVert\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2}\right\rVert_{W^{2,1}(\mathbb{B}^{d})}<\delta$ holds (the weight $|x|^{M-(d-2)}$ can be discarded since M>d2M>d-2). This proves that the conditions to apply Theorem˜4.3 are met. Inserting (4.10) and the global estimate of Lemma˜4.1 in (4.5) yields
$\left\lVert\gamma_{1}-\gamma_{2}\right\rVert_{H^{1}(\mathbb{B}^{d})}+\left\lVert\gamma_{1}-\gamma_{2}\right\rVert_{W^{2,1}(\mathbb{B}^{d})}<C_{N,K,d,p}\,\left\lVert\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2}\right\rVert_{W^{2,1}(\mathbb{B}^{d})}^{\alpha\beta}+|a_{1}-a_{2}|,$
since the weight $|x|^{M-(d-2)}$ can be removed from Lemma˜4.1 using M>d2M>d-2 and that $\gamma^{\mathrm{B}}\in L^{1}(\mathbb{B}^{d})$. The Hölder exponent of this estimate is just the product of the one of Theorem˜4.3, with the one of Lemma˜4.1. Finally, one can use (4.8) to bound $|a_{1}-a_{2}|$, and that, for all $1<q<p$,
@@
(A.1) $\left\lVert V_{j}\right\rVert_{L^{p}(\mathbb{B}^{d})}\leq C_{K}N.$
@@
$(r^{d-1}w^{\prime})^{\prime}=r^{d-1}g,\qquad g:=-q_{1}w-(q_{1}-q_{2})u_{2}.$
@@
since $w^{\prime\prime}(r)=\frac{d-1}{r}w^{\prime}-g$. a As a consequence, we have

### v2

@@ 3.5. Explicit examples
Here we present a family of conductivities for which the Born approximation can be computed explicitly.
@@
    The presence of the zero resonance for $d\geq 3$ results in a Born approximation $\gamma^{\mathrm{B}}$ with a logarithmic singularity at the origin (and $V^{\mathrm{B}}$ with a $|x|^{-2}$ singularity), as predicted by ˜5 (see Fig.˜1).
@@
3.6. $V^{\mathrm{B}}$ and spectral asymptotics of $\Lambda^{\mathrm{S}}_{V}$
To conclude this section, we prove a result analogous to ˜3 in the context of DtN associated to Schrödinger operators.
Theorem 3.12.
Let $d\geq 2$, $p>d/2$ and choose $s\in(0,1)$. For any two potentials $V_{1},V_{2}\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$ such that $0\notin\operatorname{\mathrm{Spec}}_{H^{1}_{0}(\mathbb{B}^{d})}(-\Delta+V_{j})$, $j=1,2$, the following are equivalent.
  1. i)
    $V_{1}|_{U_{s}}=V_{2}|_{U_{s}}$.
  2. ii)
    The following spectral asymptotics hold for $\Lambda^{\mathrm{S}}_{V_{1}}-\Lambda^{\mathrm{S}}_{V_{2}}$:
    $\lambda^{\mathrm{S}}_{k}[V_{1}]-\lambda^{\mathrm{S}}_{k}[V_{2}]=\mathcal{O}(s^{2k}),\qquad k\to\infty.$
The proof is based on Lemma˜2.4 and results from [DMMN24]. In fact, following ideas from that paper, one could remove the assumption $0\notin\operatorname{\mathrm{Spec}}_{H^{1}_{0}(\mathbb{B}^{d})}(-\Delta+V_{j})$ and assume weaker regularity conditions than $V_{j}\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$.
Proof.
Start by noting that (3.4) implies that
(3.14) $\lambda^{\mathrm{S}}_{k}[V_{1}]-\lambda^{\mathrm{S}}_{k}[V_{2}]=\sigma_{k}[V^{\mathrm{B}}_{1}-V^{\mathrm{B}}_{2}],\qquad k>k_{0}:=\max(k_{V_{1}},k_{V_{2}}).$
Then, if
$F(x):=(V^{\mathrm{B}}_{1}(x)-V^{\mathrm{B}}_{2}(x))|x|^{2k_{0}},$
one has that $F\in L^{1}_{\mathrm{rad}}(\mathbb{B}^{d})$. If assumption i) holds, then by [DMMN24, Theorem 2] $V^{\mathrm{B}}_{1}|_{U_{s}}=V^{\mathrm{B}}_{2}|_{U_{s}}$, which yields $F|_{U_{s}}=0$. Then, Lemma˜2.4 implies that $C>0$ exists such that, for every $k\geq k_{0}$
$|\sigma_{k-k_{0}}[F]|\leq Cs^{2(k-k_{0})},$
which suffices to conclude ii) after noticing that $\sigma_{k-k_{0}}[F]=\sigma_{k}[V^{\mathrm{B}}_{1}-V^{\mathrm{B}}_{2}]$.
To prove the converse, note that ii) implies, in view of (3.14) and (3.6), that $\sigma_{k}[F]=\mathcal{O}(s^{2k})$ as $k\to\infty$. Then, Lemma˜2.4 implies $F|_{U_{s}}=0$. As a consequence, $V^{\mathrm{B}}_{1}|_{U_{s}}=V^{\mathrm{B}}_{2}|_{U_{s}}$ a.e. and [DMMN24, Theorem 2] implies that $V_{1}|_{U_{s}}=V_{2}|_{U_{s}}$ as claimed. ∎
@@ 4.1. Hölder stability
provided that $\gamma_{1},\gamma_{2}\in\mathcal{K}\cap W^{2,p}(\mathbb{B}^{d})$, for some $p>d/2$, and that $\gamma_{1}|_{\mathbb{S}^{d-1}}=\gamma_{2}|_{\mathbb{S}^{d-1}}$. Indeed, by ˜1 and (2.7) we have
$|\lambda_{k}[\gamma_{1}]-\lambda_{k}$ $[\gamma_{2}]|\leq\frac{1}{2}\left|\sigma_{k}[\Delta(\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2})]\right|+\frac{1}{2}\left|\partial_{\nu}\gamma_{1}|_{\mathbb{S}^{d-1}}-\partial_{\nu}\gamma_{2}|_{\mathbb{S}^{d-1}}\right|$
$=\frac{1}{2}\left|\sigma_{k}[\Delta(\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2})]\right|+\frac{1}{2}\left|\sigma_{0}[\Delta(\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2})]\right|$
$\leq\sup_{k\in\mathbb{N}_{0}}\left|\sigma_{k}[\Delta(\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2})]\right|\leq\left\lVert\Delta(\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2})\right\rVert_{L^{1}(\mathbb{B}^{d})}\leq\left\lVert\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2}\right\rVert_{W^{2,1}(\mathbb{B}^{d})}.$
It is well known that in general $\omega_{\mathcal{K}}$ in (4.1) cannot be taken better than logarithmic when $\mathcal{K}$ includes a ball in a Sobolev space, [Man01, KRS21]. ˜4 improves estimate (4.2) to a Hölder modulus of continuity, and gets rid of the assumption $\gamma_{1}|_{\mathbb{S}^{d-1}}=\gamma_{2}|_{\mathbb{S}^{d-1}}$.
@@ 4.2. Proof of ˜4
then, if $a_{j}:=\gamma_{j}|_{\mathbb{S}^{d-1}}$, $b_{j}:=\frac{1}{2}\partial_{\nu}\gamma_{j}|_{\mathbb{S}^{d-1}}$ one has the local estimate
@@
(4.5) $\left\lVert V_{1}-V_{2}\right\rVert_{L^{1}(U_{s})}<C_{d,N,p,s}\,\left\lVert V^{\mathrm{B}}_{1}-V^{\mathrm{B}}_{2}\right\rVert_{L^{1}(U_{s})}^{\frac{p-1}{2p-1}}.$
@@
Let $d\geq 2$, M>d2M>d-2, and $d/2<p<\infty$. There exists a constant $\beta(d,p)>0$ such that if $M\geq\beta(d,p)\left\lVert V\right\rVert_{L^{p}(\mathbb{B}^{d})}$, then
$\int_{\mathbb{B}^{d}}|V^{\mathrm{B}}(x)||x|^{M-(d-2)}\,dx\leq 2|\mathbb{S}^{d-1}|\beta(d,p)\left\lVert V\right\rVert_{L^{p}(\mathbb{B}^{d})}.$
@@
Since the conductivities are radial and (1.5) holds, by Lemma˜2.2 i) we have for some $C>0$ that
@@
where $a_{j}:=\gamma_{j}|_{\mathbb{S}^{d-1}}$, $b_{j}:=\frac{1}{2}\partial_{\nu}\gamma_{j}|_{\mathbb{S}^{d-1}}$. Locally, the same bound holds:
@@
Conditions (1.10) on $\gamma_{j}$, $j=1,2$ imply a corresponding bound for $\left\lVert V_{j}\right\rVert_{L^{p}}$ using that $\Delta\sqrt{\gamma_{j}}=V_{j}\sqrt{\gamma_{j}}$, so the first estimates in (4.4) and (4.6) hold. We now consider the global and local cases separately.
Proof of local estimate i). First we claim that for any given $0<\varepsilon<1$, there exists a $0<\delta<1$ (depending on $\varepsilon,d,K,N,s$) such that $\left\lVert V^{\mathrm{B}}_{1}-V^{\mathrm{B}}_{2}\right\rVert_{L^{1}(U_{s})}<\varepsilon$ holds if $\left\lVert\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2}\right\rVert_{W^{2,1}(\mathbb{B}^{d})}<\delta$ holds. This can be proved using (2.15) together with $K^{-1}\leq\gamma_{j}(x)\leq K$ which imply that
$\left\lVert V^{\mathrm{B}}_{1}-V^{\mathrm{B}}_{2}\right\rVert_{L^{1}(U_{s})}=\left\lVert(2a_{1})^{-1}(2a_{1}V^{\mathrm{B}}_{1}-2a_{2}V^{\mathrm{B}}_{2})+a_{2}V^{\mathrm{B}}_{2}(a_{1}^{-1}-a_{2}^{-1})\right\rVert_{L^{1}(U_{s})}\\ \leq\frac{K}{2}\left\lVert\Delta\gamma^{\mathrm{B}}_{1}-\Delta\gamma^{\mathrm{B}}_{2}\right\rVert_{L^{1}(U_{s})}+|a_{1}^{-1}-a_{2}^{-1}||a_{2}|\left\lVert V^{\mathrm{B}}_{2}\right\rVert_{L^{1}(U_{s})},$
we now use that $|a_{1}^{-1}-a_{2}^{-1}|\leq C_{K}|a_{1}-a_{2}|$, since $K^{-1}\leq a_{j}\leq K$, Lemma˜4.4 to bound the norm of $V^{\mathrm{B}}_{2}$ (the weight in the lemma is harmless since we are in $U_{s}$), and (4.9), to obtain
@@
$\left\lVert\gamma_{1}-\gamma_{2}\right\rVert_{W^{2,1}(U_{s})}<C_{N,K,d,p,s}\,\left\lVert\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2}\right\rVert_{W^{2,1}(U_{s})}^{\frac{p-1}{2p-1}}+|a_{1}-a_{2}|+|b_{1}-b_{2}|.$
@@
Proof of global estimate ii). Let M>d1M>d-1. As in the local case we have
$\int_{\mathbb{B}^{d}}|V^{\mathrm{B}}_{1}(x)-V^{\mathrm{B}}_{2}(x)||x|^{M-(d-2)}\,dx\leq\frac{K}{2}\int_{\mathbb{B}^{d}}|\Delta\gamma^{\mathrm{B}}_{1}(x)-\Delta\gamma^{\mathrm{B}}_{2}(x)||x|^{M-(d-2)}\,dx\\ +|a_{1}^{-1}-a_{2}^{-1}||a_{2}|\int_{\mathbb{B}^{d}}|V^{\mathrm{B}}_{2}(x)||x|^{M-(d-2)}\,dx,$
By Proposition˜2.1 iv) and (2.15), $\Delta\gamma^{\mathrm{B}}_{j}\in L^{1}(\mathbb{B}^{d})$ for $j=1,2$. Hence $|x|^{M-(d-2)}$ can be removed in the first integral in the RHS since M>d2M>d-2. Using Lemma˜4.4, (1.10), (4.8) and the same arguments used to prove (4.10) we obtain
(4.11) $\int_{\mathbb{B}^{d}}|V^{\mathrm{B}}_{1}(x)-V^{\mathrm{B}}_{2}(x)||x|^{M-(d-2)}\,dx\leq C(d,K,N,M)\left\lVert\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2}\right\rVert_{W^{2,1}(\mathbb{B}^{d})},$
for $M$ large enough. We can now fix the value of $M$ (the size of $M$ depends on the bounds on $\left\lVert V_{j}\right\rVert_{L^{p}(\mathbb{B}^{d})}$, and hence on $N$ that controls the size of conductivities). Therefore, there exists a $\delta$ such that, if $\left\lVert\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2}\right\rVert_{W^{2,1}(\mathbb{B}^{d})}<\delta$ holds, then (4.6) holds. This proves that the conditions to apply Theorem˜4.3 are met. Inserting (4.11) and the global estimate of Lemma˜4.1 in (4.7) yields
$\left\lVert\gamma_{1}-\gamma_{2}\right\rVert_{H^{1}(\mathbb{B}^{d})}+\left\lVert\gamma_{1}-\gamma_{2}\right\rVert_{W^{2,1}(\mathbb{B}^{d})}<C_{N,K,d,p}\,\left\lVert\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2}\right\rVert_{W^{2,1}(\mathbb{B}^{d})}^{\alpha\beta}+|a_{1}-a_{2}|.$
The Hölder exponent of this estimate is just the product of the one of Theorem˜4.3, with the one of Lemma˜4.1. Finally, one can use (4.8) to bound $|a_{1}-a_{2}|$, and that, for all $1<q<p$,
@@
$\left\lVert V_{j}\right\rVert_{L^{p}(\mathbb{B}^{d})}\leq C_{K}N.$
@@
$(r^{d-1}w^{\prime})^{\prime}=r^{d-1}g,\qquad g:=q_{1}w+(q_{1}-q_{2})u_{2}.$
@@
since $w^{\prime\prime}(r)=\frac{d-1}{r}w^{\prime}-g$. As a consequence, we have